OpenStudy (anonymous):

Identify the vertex and the axis of symmetry for the graph y=3(x-3)^2+4

OpenStudy (anonymous):

vertex is (3,4)

OpenStudy (anonymous):

this is in the form \[y=a(x-h)^2+k\] which is a parabola with vertex (h,k) so you just read off your answer

OpenStudy (anonymous):

Would x=-3 or positive ?

OpenStudy (anonymous):

no x = 3

OpenStudy (anonymous):

reason as follows. if x = 3 then x-3=3-3 =0 so the square term is 0

OpenStudy (anonymous):

and since the square term is a SQUARE the smallest it can be is zero. so the smallest y can be is 4, and it if 4 if x is 3